# Injecting Measurement Information Yields a Fast and Noise-Robust Diffusion-Based Inverse Problem Solver

### v1

@@ D.2 Comparison against other Jacobian-Free Methods (Table 5)
A major advantage of DCS is the fact that it is Jacobian-free (Section 3.4) — this results in at least $6\times$ reduction in memory cost during inference compared with Jacobian-based methods, which can be a major enabling factor for the adoption of such algorithms on consumer GPUs and edge devices. However, naively removing the backpropagation through the score network can reduce the quality of the measurement consistency correction step in inverse solvers. In this experiment, we demonstrate that our treatment via the maximum likelihood framework and the noise-aware maximization results in significantly higher quality samples, compared to a naive implementation in DPS-JF and LGD-JF, which are both Jacobian-free variants of the original algorithms (Chung et al. 2022a) and (Song et al. 2023a). Namely, we approximate the Jacobian with respect to the input to the denoising network (left hand side) by the Jacobian with respect to the predicted $\mathbf{x}_{0}$ (right hand side)
@@ D.4 Subset of FFHQ used in other works
Table 7: Quantitative evaluation of our method on FFHQ 256x256, following the experimental setup of (Song et al. 2024). We compare against pixel-based solvers (upper half) and latent-based solvers (lower half).
@@ E.1 Our Method
Our proposed DCS has just two primary hyperparameters, as described in the table below. First is the number of time steps $T$. This has relatively little effect on our model performance on most tasks. However, it is occasionally helpful to increase $T$, especially in box inpainting, where there is zero signal from $\mathbf{y}$ in the masked region. Here, higher $T$ allows the diffusion model to obtain a better solution in this unconditional diffusion process. Second, we have the choice of minimizer, which is by default the Adam optimizer (Kingma and Ba 2014). However, in the case of linear $\mathcal{A}$, this optimizer can be replaced by the closed form analytical solution to $\mathcal{A}(\mathbf{x})=\mathbf{y}$.
@@ E.2 Latent Models on ImageNet
We note that previous latent models use the pretrained weights in (Rombach et al. 2022) for $256\times 256$ resolution datasets. However, there are no published weights in the GitHub repository for unconditional ImageNet, making a fair comparison of our method against latent models more involved. To this end, we leverage a significantly more powerful Stable Diffusion v1.5 model, with publicly available weights on HuggingFace for our experiments. The measurements and the output images are appropriately scaled for a fair comparison.
@@ E.3 STSL
At the time of writing this work, we did not find publicly available code for STSL (Rout et al. 2024). Therefore, we implement the algorithm ourselves in our codebase, and use the hyperparameters provided in the paper.
@@ E.4 ReSample
We directly use the published code of ReSample (Song et al. 2024) with no changes in our paper. We discuss two notable aspects of the experiments with ReSample. First, the implementation on GitHub differs from that pseudocode discussed in the paper. Namely, the pseudocode in the paper describes enforcing latent- and pixel-based consistency occasionally during an otherwise unconditional sampling process.
In the code we observed that the sampling step taken is actually a DPS (Chung et al. 2022a) sampling step, which includes a posterior-based guidance step that takes an expensive gradient of the noise function. To see this, note that L255 in the resample_sampling function in ddim.py calls a function measurement_cond_fn, which is defined at L62 in main.py and passed into the resampling function. This function is a member of the class PosteriorSampling defined in L53 in condition_methods.py. Inspecting this class, we note that it calls torch.autograd.grad on the diffusion step as a function of x_prev (L33 or L39). In other words, a gradient is computed for the measurement norm with respect to the input to the diffusion model, i.e., a DPS step.
@@
In experiments, we note that ReSample is significantly slower than other algorithms during sampling (see Table 1). For example, sampling $\sim 1000$ images with ImageNet takes more than two weeks on an A6000 GPU. Since we run five different experimental conditions for each dataset, this was an unacceptably long runtime for our academic resources. Therefore, we reduce the number of diffusion steps $T$ of ReSample in our experiments, from $500$ reported in (Song et al. 2024) to $50$. However, we do provide a single experiment from the (Song et al. 2024) paper, where we reproduce the hyperparameters and dataset (a $100$ image subset of FFHQ). We note that (Song et al. 2024) took a subset of the FFHQ dataset, where performance differed from the full $\times256$-1K dataset performance (c.f. Table 2). Since the subset was not published, we selected a dataset based where ReSample obtained the same performance with its default parameters in (Song et al. 2024) (Table 7).
@@ E.5 DDRM
We used the version of DDRM which is implemented in the DDNM codebase. While DDRM may theoretically be able to handle deblurring tasks, due to the high rank of the forward operators, the SVD cannot be explicitly defined in memory, and no existing code-base for DDRM supplies fast and memory-saving versions of these operators. Because in our settings DDRM’s performance was similar or worse than that of DDNM, and due to the fact that DDRM can be considered a subtype of DDNM (see Appendix of (Wang, Yu, and Zhang 2022)), we do not run benchmarks.

### v2

@@ D.2 Comparison against other Jacobian-Free Methods (Table 5)
A major advantage of DCS is the fact that it is Jacobian-free (Section 3.4) — this results in at least $6\times$ reduction in memory cost during inference compared with Jacobian-based methods, which can be a major enabling factor for the adoption of such algorithms on consumer GPUs and edge devices. However, naively removing the backpropagation through the score network can reduce the quality of the measurement consistency correction step in inverse solvers. In this experiment, we demonstrate that our treatment via the maximum likelihood framework and the noise-aware maximization results in significantly higher quality samples, compared to a naive implementation in DPS-JF and LGD-JF, which are both Jacobian-free variants of the original algorithms Chung et al., 2022a and Song et al., 2023a . Namely, we approximate the Jacobian with respect to the input to the denoising network (left hand side) by the Jacobian with respect to the predicted $\mathbf{x}_{0}$ (right hand side)
@@ D.4 Subset of FFHQ used in other works
Table 7: Quantitative evaluation of our method on FFHQ 256x256, following the experimental setup of (Song et al.,, 2024). We compare against pixel-based solvers (upper half) and latent-based solvers (lower half).
@@ E.1 Our Method
Our proposed DCS has just two primary hyperparameters, as described in the table below. First is the number of time steps $T$. This has relatively little effect on our model performance on most tasks. However, it is occasionally helpful to increase $T$, especially in box inpainting, where there is zero signal from $\mathbf{y}$ in the masked region. Here, higher $T$ allows the diffusion model to obtain a better solution in this unconditional diffusion process. Second, we have the choice of minimizer, which is by default the Adam optimizer Kingma and Ba, (2014). However, in the case of linear $\mathcal{A}$, this optimizer can be replaced by the closed form analytical solution to $\mathcal{A}(\mathbf{x})=\mathbf{y}$.
@@ E.2 Latent Models on ImageNet
We note that previous latent models use the pretrained weights in (Rombach et al.,, 2022) for $256\times 256$ resolution datasets. However, there are no published weights in the GitHub repository for unconditional ImageNet, making a fair comparison of our method against latent models more involved. To this end, we leverage a significantly more powerful Stable Diffusion v1.5 model, with publicly available weights on HuggingFace for our experiments. The measurements and the output images are appropriately scaled for a fair comparison.
@@ E.3 STSL
At the time of writing this work, we did not find publicly available code for STSL (Rout et al.,, 2024). Therefore, we implement the algorithm ourselves in our codebase, and use the hyperparameters provided in the paper.
@@ E.4 ReSample
We directly use the published code of ReSample (Song et al.,, 2024) with no changes in our paper. We discuss two notable aspects of the experiments with ReSample. First, the implementation on GitHub differs from that pseudocode discussed in the paper. Namely, the pseudocode in the paper describes enforcing latent- and pixel-based consistency occasionally during an otherwise unconditional sampling process.
In the code we observed that the sampling step taken is actually a DPS (Chung et al., 2022a, ) sampling step, which includes a posterior-based guidance step that takes an expensive gradient of the noise function. To see this, note that L255 in the resample_sampling function in ddim.py calls a function measurement_cond_fn, which is defined at L62 in main.py and passed into the resampling function. This function is a member of the class PosteriorSampling defined in L53 in condition_methods.py. Inspecting this class, we note that it calls torch.autograd.grad on the diffusion step as a function of x_prev (L33 or L39). In other words, a gradient is computed for the measurement norm with respect to the input to the diffusion model, i.e., a DPS step.
@@
In experiments, we note that ReSample is significantly slower than other algorithms during sampling (see Table 1). For example, sampling $\sim 1000$ images with ImageNet takes more than two weeks on an A6000 GPU. Since we run five different experimental conditions for each dataset, this was an unacceptably long runtime for our academic resources. Therefore, we reduce the number of diffusion steps $T$ of ReSample in our experiments, from $500$ reported in (Song et al.,, 2024) to $50$. However, we do provide a single experiment from the (Song et al.,, 2024) paper, where we reproduce the hyperparameters and dataset (a $100$ image subset of FFHQ). We note that (Song et al.,, 2024) took a subset of the FFHQ dataset, where performance differed from the full $\times256$-1K dataset performance (c.f. Table 2). Since the subset was not published, we selected a dataset based where ReSample obtained the same performance with its default parameters in (Song et al.,, 2024) (Table 7).
@@ E.5 DDRM
We used the version of DDRM which is implemented in the DDNM codebase. While DDRM may theoretically be able to handle deblurring tasks, due to the high rank of the forward operators, the SVD cannot be explicitly defined in memory, and no existing code-base for DDRM supplies fast and memory-saving versions of these operators. Because in our settings DDRM’s performance was similar or worse than that of DDNM, and due to the fact that DDRM can be considered a subtype of DDNM (see Appendix of Wang et al., (2022)), we do not run benchmarks.